OpenStudy (anonymous):

solve for x: log3(2x-5)=log3(x+4)

OpenStudy (anonymous):

x=9

OpenStudy (anonymous):

See when you have: \[\large \log_x(a) = \log_x(b)\] This implies : \[\large \color{green}{a = b}\] So apply this in your question, you will get: \[\large 2x- 5 = x + 4\] Solve for x here.. 1. Add 5 both the sides first. 2. Subtract x from both the sides, and you will get your x..

OpenStudy (anonymous):

log3(2x-5)=log3(x+4) log(2x-5)/log3 = log(x+4)/log3 log(2x-5)=log(x+4) appyling antilog antilog(log(2x-5))=antilog(log(x+4)) 2x-5 = x+4 2x-x = 4+5 x=9(ans..)

OpenStudy (anonymous):

Welcome to OpenStudy @M16gogirl